How much energy is in the elastic potential energy store of a spring with a spring constant of 4 N/m if it is stretched from 1 m to 1.1 m in length?


1
Expert's answer
2022-05-24T13:56:33-0400
"W=\frac{kx_2^2}{2}-\frac{kx_1^2}{2}\\
=\frac{4}{2}*(1.1^2-1.0^2)=0.42\:\rm J"